# Synthesizing High-Quality Programming Tasks with LLM-based Expert and Student Agents

### v1

@@ 6.2 Real-world Performance of Our Web Application
Statistic Participant ID Average
P1 P2 P3 P4 P5 P6 P7 P8 P9 P10
No. Requests $6$ $5$ $6$ $6$ $5$ $5$ $6$ $5$ $5$ $5$ $5.40$
No. Synthesized Tasks $5$ $5$ $5$ $5$ $5$ $5$ $5$ $5$ $5$ $5$ $5.00$
No. Solved Tasks $5$ $5$ $4$ $4$ $4$ $5$ $5$ $5$ $3$ $5$ $4.50$
Figure 9: Statistics of participants when requesting and solving tasks synthesized by PyTaskSyn on our web application. Our system successfully synthesized tasks for 92.6% of participant requests.
@@ 7 Concluding Discussions
In this work, we introduced PyTaskSyn, a novel synthesis technique that leverages generative models as agents simulating different classroom roles to validate generated programming tasks. Each stage in our pipeline contributes uniquely to the validation process, collectively ensuring the creation of tasks that are of high-quality and meet diverse learning objectives. Through extensive expert evaluation and user studies, we demonstrated that our approach significantly improves the quality of generated programming tasks with minimal cost, while maintaining reasonable coverage across various themes and programming concepts.
Our work brings two important implications for leveraging generative AI for computing education. First, our results demonstrate how we can effectively automate the quality assessment of generated programming tasks. This paves the way to reducing educator workload when designing practice exercises that target specific learning objectives in programming education. Second, our results show that by breaking down the task synthesis process into different stages, we can leverage generative models that simulate different agents for specialized roles. This agent-based approach opens up new opportunities for utilizing generative models to simulate learning analytics to generate and validate educational content.
Next, we discuss some limitations of our work and possible ways of approaching them in the future. First, we adopt an accept/reject approach during our validation; future work could employ a framework for refining programming tasks by leveraging feedback from the validation stage. Second, we do not analyze whether the passing threshold of simulated students in our pipeline affects the synthesized task’s difficulty; it would be valuable to investigate it and see whether it aligns with difficulty assessed by experts. Third, we focus on Python programming; it would be interesting to explore whether the capabilities of generative models for validating tasks extend beyond Python. Fourth, we conducted studies with a relatively small number of participants; it would be important to conduct larger-scale studies with students and assess their learning outcomes.

### v2

@@ 7 Concluding Discussions
We introduced PyTaskSyn, a novel synthesis technique that leverages generative models as agents simulating different classroom roles to validate generated programming tasks. Each stage in our pipeline contributes uniquely to the validation process, collectively ensuring the creation of tasks that are of high-quality and meet diverse learning objectives. Through extensive expert evaluation and user studies, we demonstrated that our approach significantly improves the quality of generated programming tasks with minimal cost, while maintaining reasonable coverage across various themes and programming concepts.
Our work brings two important implications for leveraging generative AI for computing education. First, we demonstrate how to effectively automate the quality assessment of generated programming tasks. This paves the way to reducing educator workload when designing practice exercises that target specific learning objectives in programming education. Second, our results show that by breaking down the task synthesis process into different stages, we can leverage generative models that simulate different agents for specialized roles. This agent-based approach opens up new opportunities for utilizing generative models to simulate learning analytics to generate and validate educational content.
Next, we discuss some limitations of our work and possible ways of approaching them in the future. First, we adopt an accept/reject approach during our validation; future work could employ a framework for refining programming tasks by leveraging feedback from the validation stage. Second, we do not analyze whether the passing threshold of simulated students affects the synthesized task’s difficulty; it would be valuable to investigate it and see whether it aligns with difficulty assessed by experts. Third, we focus on Python programming; it would be interesting to explore whether the capabilities of generative models for validating tasks extend beyond Python. Fourth, we conducted studies with a relatively small number of participants; it would be important to conduct larger-scale studies with students and assess their learning outcomes.